OpenStudy (anonymous):

52. Solve ln2 + lnx = 5. Round to the nearest thousandth, if necessary

OpenStudy (marigirl):

first subtract ln(2) from both sides

OpenStudy (marigirl):

ln (x)= 5- ln(2) which is the same as ln(x)=5+ln(0.5) now we can isolate the x on the left side by log e to both sides x=e^(5+ln(0.5)) then the e and ln will cancel out x=e^(5)*0.5